OpenStudy (anonymous):

why is the slope of 1/x-1 at x=2 1?

OpenStudy (anonymous):

It is a horizontal translation of 1/x, shifted to the right one unit. 1/x has a slope of 1 at x = 1 so this function has that slope at x = 2.

OpenStudy (anonymous):

Or you could try the definition of a derivative...

OpenStudy (anonymous):

\[\frac{ 1 }{ x-1 } \] at x = 2, so it's translated down 1 unit.

OpenStudy (anonymous):

i'm trying to find the slope of the curve

OpenStudy (phi):

do you know how to find the derivative? if not, use \[ \frac{f(x+h)-f(x)}{h} \] and let h->0 to find the slope. then plug in x=2 to get the value of the slope

OpenStudy (anonymous):

\[ \frac{f(\frac{ 1 }{ 2-1 }) +h - f(2) }{ h }\] \[\frac{ 1+h-1}{ h} = 1\]

OpenStudy (anonymous):

i do know how to do the derivative... i was really unsure lol

OpenStudy (phi):

in this case \[ f(x)= \frac{1}{x-1} \] and \[ f(x+h)= \frac{1}{(x+h)-1}\] can you finish?

OpenStudy (phi):

\[\lim_{h \rightarrow 0}\frac{ 1 }{ h}\left( \frac{ 1 }{ x+h-1 } -\frac{ 1 }{ x-1 } \right)\]

OpenStudy (anonymous):

well i got one either way. is my way correct or wrong?

OpenStudy (phi):

But the answer is not 1. to finish \[ \lim_{h \rightarrow 0}\frac{ 1 }{ h}\left( \frac{ x-1 }{ (x-1)(x+h-1) } -\frac{ x+h-1 }{ (x-1)(x+h-1) } \right)\] \[ \lim_{h \rightarrow 0}\frac{ 1 }{ h}\left( \frac{ -h }{ (x-1)(x+h-1) } \right)\] \[ \lim_{h \rightarrow 0}\left( \frac{ -1 }{ (x-1)(x+h-1) } \right)\] now let h go to zero to get \[ \frac{ -1 }{ (x-1)(x-1) } \] or \[ f'(x)= \frac{-1}{(x-1)^2} \] plug in x=2 to find the slope at x=2 you get -1

OpenStudy (phi):

once you learn derivatives \[ f(x)= (x-1)^{-1} \] and \[ f'(x)= -(x-1)^{-2} \] follows immediately from the rule on differentiating powers.